a) if three samples of KHP weighing , 31.21 mg +- 4.3 % ; 81.4 mg +- 7.12% and 213.2 mg +- 8.4% were added together, determine the total mass of KHP and its absolute uncertainty. Express the anwer with the correct number of significant figures.

b) If the mass of KHP from part(a) of this problem were dissolved in 207.1 ml +- 0.9 ml of water, what would the concentration of KHP (expressed as g/ml) be ? express your answer and its absolute uncertainty with the appropriate number of significant figures.
Answer #1

Solution :-

if the masses of the KHP added together then total mass of KHP = 31.21 mg + 81.4 mg +213.2 mg = 325.8 mg

Absolute uncertainty = (31.21*4.3%/100%)+(81.4*7.12%/100%)+(213.2*8.4%/100%) = 25.0 mg

Now lets calculate the concentration of the KHP in g/ml

(325.8 mg * 1 g / 1000 mg)= 0.3258 g

0.3258 g / 207.1 ml = 0.001573 g/ml

its absolute uncertainty is = 0.001573 g/ml * 0.4345 =6.84*10^-4